Could someone please help me solve the system of inequalities
5(-x+1)+6<=6 and 5(x -1) + 6 <=6?
Thanks in advance

Answer by checkley71(8403)

5(-X+1)+6<=6
-5X+5+6<=6
-5X+11<=6
-5X<=6-11
-5X<=-5
X>=-5/-5 [NOTE: WHEN DIVIDING OR MULTIPLYING BY A NEGATIVE VALUE THE < SIGN CHANGES TO A > SIGN.
X>=1 ANSWER.
5(X-1)+6<=6
5X-5+6<=6
5X+1<=6
5X<=6-1
5X<=5
X<=5/5
X<=1 ANSWER.